If A and B are mutually exclusive events with P(A) = 0.40, then P(B):


A. can be any value between 0 and 1.

B. cannot be larger than 0.40.

C. cannot be larger than 0.60.

D. cannot be determined with the information give.

Answer by rajagopalan(174)

If A and B are mutually exclusive events
with P(A) = 0.40,
Formula P(A)+P(B)=1
then P(B): =1-P(A)
=1-0.4
=0.6
S0 the answer is C